An electron starts from rest 29.5 cm from a fixed point charge with Q = -4.00 nC How fast will the electron be moving when it is very far away? Express your answer with the appropriate units.
Answer #1

Initial potential energy will be converted to kinetic energy at far point

k*q1*q2/r^2 = 0.5*m*v^2

Here:

q1=charge of electron =-1.6*10^-19 C

q2 = -4*10^-9 C

r = 29.5 cm = 0.295 m

m=mass of electron = 9.1*10^-31 Kg

(9*10^9)*(1.6*10^-19)*(4*10^-9) / (0.295)^2 = 0.5*(9.1*10^-31)*v^2

v=1.21*10^7 m/s